A series R L circuit is connected to a 110V ac source. If the voltage across the resistor is 85 V, find the power dissipated in the circuit .


1
Expert's answer
2021-05-28T02:26:04-0400

Since the voltages across the inductor and the resistor are "90^0" out of phase, we obtain the voltage across the inductor in the phase domain as follows: If the voltage across the resistor is "V_1" , the voltage across the inductor will be "jV_2" . Assume the voltage of the source is "V_3"


Applying KVL,

"V_1+jV_2=V_1cos(\omega t)+V_2sin(\omega t)=V_3cos (\omega t-\theta^0)=85+V_2sin (\omega t)=110cos(\omega t-\theta^0)"

"110=\sqrt{85^2+V_2^2}=V_2=\sqrt{110^2-85^2}=69.82 V"